# Adversarial Agents: Black-Box Evasion Attacks with Reinforcement Learning

### v1

@@ 4 Evaluation
  1. RQ1.
    Is reinforcement learning suitable for this task, and does it improve the effectiveness and efficiency of adversarial example generation?
  2. RQ2.
    How do hyperparameters $\epsilon$ in RL Max Loss and $c$ in RL Min Norm affect adversarial example effectiveness and efficiency?
  3. RQ3.
    Do RL attacks generalize to unseen data and different victim models, and how do they compare to known black-box attacks?
@@ 4.1 Experimental Setup
We use the CIFAR-10 image dataset throughout. This dataset is comprised of $70k$ object color images separated into $60k$ training and $10k$ test examples. CIFAR-10 has been widely used to study AML in many models and training architectures. Each RGB image has $3072$ features encoding 32x32 pixel values. For black-box experiments, we use the $10k$ testing examples and create two partitions: $5k$ images for $\mathcal{D}$ in attack training, and $5k$ images as a hidden dataset $\mathcal{D}^{\prime}$ (for policy evaluation). The RL attack training was performed on NVIDIA A100 GPUs with 40 GB of VRAM on pools of the Center for High Throughput Computing (Center for High Throughput Computing, 2006).
Model Architectures. The victim model $Z$ has a ResNet50 (He et al., 2015) architecture implemented in Pytorch. It is trained on the CIFAR-10 training dataset and has a 95.67% accuracy on the testing dataset. Further training details can be found in Appendix A.
For RL training, we use the PPO implementation in StableBaselines3 (Raffin et al., 2021). The adversary’s policy consists of a ResNet18 feature extractor followed by a filly connected feedforward neural network that outputs the parameters of the action distribution. The PPO algorithm also uses a separated value function network which we model as a feedforward network that shares the policy’s ResNet18 feature extractor. PPO hyperparameters and training details are in Appendix A.
@@
Action Hyperparameters. In Section 3.1, we define the actions as selecting $N$ features of the input with associated features $\delta_{i}$ for $i\in\{1,..,N\}$. For our evaluation, we fix $N=3$ and $|\delta_{i}|\leq 0.01$ for all $i\in\{1,2,3\}$ to focus on evaluating the agents ability to learn and generalize adversarial examples.
@@
Metrics. Adversarial examples in all threat models are benchmarked on distortion and victim model misclassification. As such, we use $\ell_{2}$-norm distortion and misclassification (i.e., $\arg\max_{i}Z(x_{T})^{(i)}\neq y$) at terminal states $s_{T}$. The exception appears with black-box threat models, where the number of victim model queries per adversarial example are also considered. Thus, we track the number of victim model queries in each episode. These metrics encapsulate the strength of a black-box adversary and will be used to show their changes as the agent is trained.
@@ 4.2 Training Adversarial Agents
In this section we ask RQ1: is reinforcement learning suitable for this task, and does it improve the effectiveness and efficiency of adversarial examples? In other words, is the agent able to learn a policy with greater average return, and does the adversary get stronger with more successful adversarial examples in fewer queries? To answer this, we perform an evaluation in two parts: (1) analyzing the return and adversarial example metrics over training, and (2) policy evaluation on adversarial examples before and after training. We use average return, average $\ell_{2}$-norm distortion, misclassification rate, and median queries of adversarial examples as metrics. We observe that ($a$) if the average return increases, the agent is effectively learning a better policy on the task, ($b$) if the distortion decreases, the agent is learning more effective adversarial examples and ($c$) If median queries decrease, the agent is learning adversarial examples more efficiently.
Training Evaluation In Section 3.1, we define the rewards $R_{\text{Max Loss}}$ and $R_{\text{Min Norm}}$ for the RL Max Loss and RL Min Norm attacks, respectively. Our goal is to show that the average return for each attack increases over training to determine if the task is suitable for RL. Here, we run each attack type for $5000$ episodes against the ResNet50 victim model for 10 trials each and plot the rolling average over the previous $100$ episodes of average return, misclassification rate, median number of queries, and average $\ell_{2}$-norm distortion. We fix $c=1.0$ and $\epsilon=0.5$ for the RL Min Norm and RL Max Loss attacks, respectively. We plot the four metrics with respect to training episode with a 95% confidence interval in Figure 3.
(a) Average Return
(b) Misclassification Rate
(c) Median Queries
(d) Average $\ell_{2}$-norm Distortion
Figure 3: RL Max Loss and RL Min Norm attack training. Average return, misclassification rate, median queries, and average $\ell_{2}$-norm distortion with respect to training episodes for 10 trials per attack with a 95% confidence interval.
We observe that the average return increases over the $5000$ episodes of training in 3(a). This means that the agent is learning a better policy with more discounted reward— an artifact of RL suitability for a given task. The RL Max Loss is offset above the RL Min Norm because the latter penalizes distortion in the reward, whereas the former explicitly projects the distortion onto the specified $\epsilon$ budget.
We also see that the misclassification rate increases by 19.4% in 3(b) and the median number of queries decreases by 53.2% in 3(c), on average. This means that episodes are terminating on victim model misclassifications with higher probability. Further, the agent gets to the misclassifications more efficiently because the median number of queries reduces from beginning to end of training. We see the difference in attack performance in 3(d) where RL Max Loss averages 0.37 $\ell_{2}$-norm distortion throughout training and RL Min Norm increases logarithmically to 0.48 $\ell_{2}$-norm distortion with high variance. This is because the distortion budget is enforced in the former and softly constrained in the latter. Consequently, the misclassification rate in RL Min Norm is 8.3% higher at the end of training. Nevertheless, the results demonstrate RL’s ability to learn adversarial examples.
Policy Evaluation With our black-box evaluation metrics for adversarial examples, we aim to show that adversarial examples produced from a trained policy are more efficient than those produced by a random policy. Here, we take 10 random policies (Random) and 10 trained policies (Trained) for each RL attack after $5000$ episodes of training and evaluate each on 100 images randomly sampled from $\mathcal{D}$. This equates to 1000 episodes of policy evaluation for each RL attack.
In RL Max Loss, the trained agent misclassification rate in the test setting is 75.0% compared to the random agent with 59.1%. Similarly, the RL Min Norm trained agent achieves 86.6% misclassification rate compared to the random agent with 60.7%. We plot the distribution of queries per adversarial example in Figure 4 normalized over the total number of adversarial examples to better understand if the adversarial examples are not only getting more successful but doing so efficiently.
In both attacks, we observe that the adversarial examples crafted with the trained policy are heavily right skewed compared to the random policy. In the RL Max Loss attack distribution (4(a)), we note that the median number of queries for Trained and Random policies are 554 and 1001, respectively. Similarly, in the RL Min Norm attack distribution (4(b)), the median number of queries for Trained and Random policies are 608 and 961, respectively. The steep shift in query distribution of adversarial examples from before training to after training tells us that the RL agents learn an efficient attack strategy to generate more adversarial examples.
(a) RL Max Loss
(b) RL Min Norm
Figure 4: RL Max Loss and RL Min Norm adversarial example query distribution normalized by misclassification rate for trained agents (Trained) and random agents (Random) in each RL attack setting.
4.3 RL Attack Hyperparameter Sensitivity
@@
In Section 3.1, we define two types of RL attacks with shared state and action spaces but different reward and transition functions. Recall that the $\epsilon$ parameter in RL Max Loss controls the maximum distortion allowable, and $c$ in RL Min Norm controls the weight of the reward given to reducing victim model confidence over minimizing distortion. Here, we target RQ2: how do hyperparameters $\epsilon$ in RL Max Loss and $c$ in RL Min Norm affect adversarial example effectiveness and efficiency?
@@
This analysis is broken down into two parts: (1) analyzing the effect $\epsilon$ has on the adversarial examples crafted by RL Max Loss, and (2) analyzing the effect $c$ has on the adversarial examples crafted by RL Min Norm. We use the misclassification rate and $\ell_{2}$-norm distortion on adversarial examples crafted after training to examine how changing $\epsilon$ and $c$ affects adversarial capabilities. We plot the performance of a random agent and trained agent in a policy evaluation setting in Figure 5.
@@
Analyzing the effect of $\epsilon$. With our setup of the RL Max Loss attack, the agent is rewarded by $\Delta_{t+1}f$ at each timestep and all distortions stay within $\epsilon$ $\ell_{2}$-norm distortion budget through $R_{\text{Max Loss}}$ and $P_{\text{Min Norm}}$. We run the RL Max Loss attack for $5000$ episodes with $\epsilon$ from 00 to $1.0$ in $0.1$ increments against the ResNet50 victim model and each tested on 100 images randomly sampled from $\mathcal{D}$. We plot the random and a trained agent misclassification rate with respect to $\epsilon$ in 5(a).
@@
An adversary’s misclassification rate will increase with respect to distortion budget because the adversary is given more space to maneuver. Indeed, we observe that the trained RL Max Loss agent’s performance increases logarithmically with respect to $\epsilon$. We note that because the adversary learns a better policy, the trained agent’s performance tops the random agent by an average of 13.2% across $\epsilon$ values. Depending on the threat model and the adversary’s distortion budget goal, we can observe the performance/budget trade-off when selecting $\epsilon$ to train the agent.
@@
Analyzing the effect of $c$. In our setup of the RL Min Norm attack, the agent is rewarded by $c\cdot\Delta_{t+1}f+\Delta_{t+1}\delta$ at each timestep through $R_{\text{Min Norm}}$ and $P_{\text{Min Norm}}$. We run the RL Min Norm attack for $5000$ episodes with $c$ from 00 to $10^{2}$ in log-scale increments against the ResNet50 victim model and each tested on 100 images randomly sampled from $\mathcal{D}$. We evaluate a random policy before training (Random) and a trained policy (Trained) for 10 trials each and plot the misclassification rate with respect to $\epsilon$ on a 95% confidence interval in 5(b). To reduce legend overload, we denote the trained agent performance as solid lines and random agent performance as dotted.
(a) RL Max Loss ($\epsilon$)
(b) RL Min Norm ($c$)
Figure 5: Hyperparameter sensitivity: (Left) RL Max Loss attack misclassification rate vs. $\epsilon$ distortion budget constraint and (Right) RL Min Norm attack misclassification rate and average $\ell_{2}$-norm distortion vs. $c$. Trained agent (Trained) results are plotted against a random agent (Random) for 10 trials each with a 95% confidence interval.
@@
The adversary’s performance and distortion will increase with respect to $c$ because the agent values decreasing the norm over decreasing victim model confidence. Indeed, we observe this behavior as $c$ increases on log-scale. Further, we see that the agent learns a policy that increases both misclassification rate and distortion. At $c=1.0$, we observe that the trained agent has a misclassification rate 29.8% higher on average and an average $\ell_{2}$-norm distortion 0.8 higher on average. These results are consistent with traditional Min Norm attacks in AML (Carlini & Wagner, 2017). Like our analysis on $\epsilon$ for RL Max Loss, an optimal value of $c$ depends on the adversary’s goals and capabilities.
4.4 RL Attack Generalization and Comparison
Each adversarial agent trains on images from the training dataset $\mathcal{D}$. To investigate a broader landscape of attack capabilities in black-box settings, we target RQ3: do RL attacks generalize to unseen data and different victim models, and how do they compare to known black-box attacks?
We bifurcate the analysis of this question into two experiments: (1) analyzing the adversarial examples crafted on an unseen dataset and (2) comparing the performance of the RL attacks against a known highly-performant and effective black-box attack, SquareAttack (Andriushchenko et al., 2020). The first part will examine adversarial examples crafted after training on randomly selected inputs from the training dataset $\mathcal{D}$ and testing dataset $\mathcal{D}^{\prime}$. The second part will compare the RL attacks against SquareAttack over misclassification rate, median queries, and average distortion.

### v2

@@ 4 Evaluation
  1. 1.
    Do adversarial agents learn more effective and efficient attacks during training?
  2. 2.
    How do key environment hyperparameters ($\epsilon$ and $c$) influence the trade-off between attack effectiveness and efficiency?
  3. 3.
    How well do trained adversarial agents generalize to unseen data, and how does it compare to traditional black-box methods?
@@ 4.1 Experimental Setup
Figure 3: RL Max Loss and RL Min Norm attack training: attack success rate (ASR), average queries on successful attacks (AQ), and average $\ell_{2}$-norm distortion on successful attacks ($\ell_{2}$) with respect to policy updates for 3 trials per attack with a 95% confidence interval on CIFAR-10 and SVHN datasets.
Model Architecture CIFAR-10 SVHN
ResNet-50 [he_deep_2015] 97.1% 96.4%
VGG-16 [simonyan_very_2015] 94.7% 95.2%
ViT-B/16 [dosovitskiy_image_2021] 97.9% 97.3%
Table 1: Test Classification Accuracy of Victim Models: comparison of victim models across the standard test split of image classification datasets.
Datasets. We use two image classification datasets in our experiments: CIFAR-10 [krizhevsky_learning_nodate] and Street View House Numbers (SVHN) [noauthor_reading_nodate]. CIFAR-10 is an object recognition dataset and is widely used in AML literature. It is made up of 60 000 training and 10 000 testing samples. Each sample has 3 072 features encoding pixel values. For black-box experiments, we use the 10 000 test samples and create two partitions: 5 000 samples for $\mathcal{D}$ in attack training, and 5 000 samples as a hidden dataset $\mathcal{D}^{\prime}$ for policy evaluation. SVHN is a digit recognition dataset collected from Google Street View images by cropping out digit sequences from house number plates. The data set consists of 73,257 samples for training and 26,032 for testing. Like CIFAR-10, each sample has 3 072 features encoding pixel values and we create two partitions of the test samples: 13 016 for $\mathcal{D}$ in attack training, and 13 016 as a hidden dataset $\mathcal{D}^{\prime}$.
Model Architectures. We experiment with three different victim models $Z$: ResNet-50 [he_deep_2015], VGG-16 [simonyan_very_2015], and ViT-B/16 [dosovitskiy_image_2021]. For our experiments, we use models pre-trained on the ImageNet-1K dataset [deng_imagenet_2009] and fine-tune them on the individual target datasets. The resulting test accuracies are reported in Table 1. For RL training, we use the PPO implementation in StableBaselines3 [raffin_stable-baselines3_2021]. The adversarial agent’s policy consists of a EfficientNet [tan_efficientnet_2020] feature extractor followed by a fully connected feedforward neural network that outputs the parameters of the action distribution. The PPO algorithm also uses a separated value function network which we model as a feedforward network that shares the policy’s EfficientNet feature extractor. For training adversarial agents, we perform our experiments on NVIDIA A100 GPUs with 40 GB of VRAM.
@@
Action Hyperparameters. Our methodology defines actions as selecting $N$ features of the input with associated features $\delta_{i}$ for $i\in\{1,..,N\}$. For our evaluation, we fix $N=5$ and $\theta=0.05$, as preliminary experiments showed these values offered a good balance between attack success and action complexity (see supplementary material for selection details). We focus on evaluating the agents ability to learn and generalize adversarial strategies.
@@
Metrics. Adversarial samples in all threat models are benchmarked on distortion and victim model misclassification. As such, we use the $\ell_{2}$ distortion and attack success rate (ASR) (i.e., $\arg\max_{i}Z(x_{T})^{(i)}\neq y$) at terminal states $s_{T}$. The exception appears with black-box threat models, where the number of victim model queries per adversarial sample are also considered. Thus, we determine the average queries to victim models on episodes that produce successful attacks (AQ). These metrics encapsulate the strength of a black-box adversary and will be used to show their changes as the agent is trained.
@@ 4.2 Training Adversarial Agents
In this section, we ask: do adversarial agents learn more effective and efficient attacks during training? In other words, is the agent able to learn a policy with greater attack success, and does the adversary get stronger with more successful adversarial samples in fewer queries? To answer this, we perform an evaluation in two parts: (1) analyzing policy learning and attack effectiveness over training, and (2) analyzing attack efficiency over training. We use ASR, AQ, and $\ell_{2}$ distortion of adversarial samples as metrics. In Figure 3, we train RL Max Loss and RL Min Norm agents in training datasets $\mathcal{D}$ for 1200 PPO policy updates with 3 random seeds each and plot the mean and 95% confidence interval over attack metrics. We fix $\epsilon=0.3$ for all RL Max Loss experiments, $c=10^{-2}$ for RL Min Norm on CIFAR-10, and $c=10^{-3}$ for RL Min Norm on SVHN. Figure 3 organizes the plots with the dataset and metrics on the columns and the RL attack type on the rows. We observe that the agent is learning (a) a more effective policy if attack success increases, (b) a more query-efficient policy if successful episode queries decrease, and (c) a more distortion-efficient policy if $\ell_{2}$ distortion decreases.
Effective Learning. Our goal is to show that the ASR for each attack increases over training to determine if the adversarial agents learn more effective attacks. Here, we focus on the ASR columns of Figure 3. Note that the ASR scales differ across RL attack type and dataset. We observe that the RL Max Loss attack consistently increases ASR over training across all models and datasets with averages of 8.7% increase on ViT-B/16, 9.3% increase on ResNet-50, and 13.2% increase on VGG-16. Similarly, we observe that the RL Min Norm attack increases the attack success with averages of 5.1% increase on ViT-B/16, 5.6% increase on ResNet-50, and 9.7% increase on VGG-16. We note that the difference in performance between RL attack types is because the RL Max Loss attack enforces the $\ell_{2}$ budget constraints in the state transition while RL Min Norm must learn it through the reward feedback. Nevertheless, the results demonstrate RL’s ability to learn effective strategies for generating adversarial samples.
Efficient Learning. With our black-box evaluation metrics for adversarial samples, we aim to show that adversarial samples produced throughout training become more efficient. Here, we focus on the AQ and $\ell_{2}$ distortion columns of Figure 3. Note that the scale of successful queries and distortion differ across RL attack type and dataset. We observe that RL Max Loss improves query-efficiency significantly throughout training with up to 11.3% and 16.9% decrease in average queries on successful episodes on CIFAR-10 and SVHN tasks, respectively. In contrast, the RL Min Norm attack improves query efficiency only on VGG-16 models by an average of 6.6%. However, these steep changes in query-efficiency for RL Max Loss tells us that the adversarial agents learn a query-efficient attack strategy.
In the $\ell_{2}$ distortion plots, we observe the same trends for each attack on both CIFAR-10 and SVHN tasks. RL Max Loss consistently stays below its $\epsilon=0.3$ budget constraint, confirming that the agent efficiently utilizes the allowable distortion to increase attack effectiveness and query-efficiency. The decrease in distortion observed in all models and datasets is influenced by the agent requiring fewer queries over training. Contrarily, the $\ell_{2}$ distortion RL Min Norm requires for attacks varies widely on the dataset and victim model. This demonstrates that the agent, while achieving marginal learning gains, struggles to efficiently minimize distortion and queries simultaneously when those constraints are only guided through the reward function, rather than being explicitly enforced in the state space like in RL Max Loss.
Figure 4: RL Max Loss Hyperparameter Sensitivity: attack success rate (ASR) versus Epsilon ($\epsilon$) for trained agents averaged over 3 random seeds.
Figure 5: RL Min Norm Hyperparameter Sensitivity: attack success rate (ASR) and $\ell_{2}$ distortion vs. $c$ for trained agents averaged over 3 random seeds.
CIFAR-10 SVHN
ResNet-50 VGG-16 ViT-B/16 ResNet-50 VGG-16 ViT-B/16
Attack & Dataset ASR AQ $\ell_{2}$ ASR AQ $\ell_{2}$ ASR AQ $\ell_{2}$ ASR AQ $\ell_{2}$ ASR AQ $\ell_{2}$ ASR AQ $\ell_{2}$
RL Max Loss ($\mathcal{D}$) 0.55 295 0.29 0.64 265 0.29 0.30 321 0.30 0.43 287 0.29 0.59 278 0.29 0.32 336 0.30
RL Max Loss ($\mathcal{D}^{\prime}$) 0.59 315 0.30 0.64 259 0.30 0.28 332 0.30 0.45 285 0.30 0.60 255 0.30 0.32 344 0.30
RL Min Norm ($\mathcal{D}$) 0.58 168 0.32 0.53 111 0.25 0.19 155 0.11 0.49 265 0.34 0.65 191 0.41 0.20 187 0.15
RL Min Norm ($\mathcal{D}^{\prime}$) 0.62 155 0.35 0.55 103 0.29 0.17 148 0.12 0.53 249 0.36 0.65 186 0.42 0.18 192 0.18
Square [andriushchenko_square_2020] ($\mathcal{D}^{\prime}$) 0.53 335 0.30 0.61 350 0.30 0.31 344 0.30 0.48 263 0.30 0.48 273 0.30 0.29 262 0.30
HSJA [chen_hopskipjumpattack_2020] ($\mathcal{D}^{\prime}$) 0.31 681 0.30 0.36 637 0.30 0.13 904 0.30 0.30 705 0.30 0.38 639 0.30 0.12 923 0.30
Bandits [ilyas_prior_2019] ($\mathcal{D}^{\prime}$) 0.44 635 0.30 0.47 662 0.30 0.08 516 0.30 0.39 556 0.30 0.46 537 0.30 0.14 589 0.30
Table 2: Performance comparison of RL-based attacks and black-box attack baselines on the CIFAR-10 and SVHN datasets across different victim models: ResNet-50, VGG-16, and ViT-B/16. We report the attack success rate (ASR), average queries on successful attacks (AQ), and average $\ell_{2}$-norm distortion on successful attacks ($\ell_{2}$) on training set $\mathcal{D}$ and testing set $\mathcal{D}^{\prime}$.
4.3 Sensitivity Analysis
@@
The two types of RL attacks share the same state and action representations, but differ in reward and transition functions. Recall that the $\epsilon$ parameter in RL Max Loss controls the maximum distortion allowable, and $c$ in RL Min Norm controls the weight of the reward given to reducing victim model confidence over minimizing distortion. Here, we ask: how do key environment hyperparameters ($\epsilon$ and $c$) influence the trade-off between attack effectiveness and efficiency?
@@
This analysis is broken down into two parts: (1) analyzing the effect $\epsilon$ has on the adversarial samples crafted by RL Max Loss, and (2) analyzing the effect $c$ has on the adversarial samples crafted by RL Min Norm. We use the ASR and $\ell_{2}$-norm distortion on adversarial samples crafted after training to examine how changing $\epsilon$ and $c$ affects adversarial capabilities. We plot the performances of trained agents in policy evaluation settings for RL Max Loss and RL Min Norm in Figure 4 and Figure 5, respectively.
@@
Analyzing the effect of $\epsilon$. With our setup of the RL Max Loss attack, the agent is rewarded by $\Delta_{t+1}f$ at each timestep and all distortions stay within $\epsilon$ $\ell_{2}$-norm distortion budget through $R_{\text{Max Loss}}$ and $P_{\text{Max Loss}}$. We run the RL Max Loss attack for 1200 PPO policy updates with $\epsilon\in\{0.1,0.3,0.5\}$ and tested on the training images from $\mathcal{D}$. We evaluate the policies trained from 3 random seeds and plot the average ASR with respect to $\epsilon$ on each victim model and dataset in Figure 4.
@@
An adversary’s ASR will increase with respect to distortion budget because the adversary is given more space to maneuver. Thus, an ASR corresponding to $\epsilon=0.0$ is a representation of one minus the victim model’s accuracy on unperturbed data. Indeed, we observe that the trained RL Max Loss agent’s performance increases dramatically with respect to $\epsilon$. Depending on the threat model and the adversary’s distortion budget goal, we can observe the performance/budget trade-off when selecting $\epsilon$ to train the agent.
@@
Analyzing the effect of $c$. In our setup of the RL Min Norm attack, the agent is rewarded by $\Delta_{t+1}f+c\cdot\Delta_{t+1}\delta$ at each timestep through $R_{\text{Min Norm}}$ and $P_{\text{Min Norm}}$. We run the RL Min Norm attack for 1200 PPO policy updates with $c\in\{10^{-3},10^{-2},10^{-1}\}$ and each tested on the training images from $\mathcal{D}$. We evaluate the policies trained from 3 random seeds and plot the average ASR (left axis) and $\ell_{2}$ distortion (right axis) with respect to $c$ on each victim model and dataset in Figure 5.
@@
The adversary’s performance and distortion will decrease with respect to $c$ because the agent values decreasing the distortion over decreasing victim model confidence. Indeed, we observe this behavior as $c$ increases on log-scale. These results are consistent with traditional Min Norm attacks in AML [carlini_towards_2017] and agents trained under the RL Min Norm approach must consider this before or during training adversarial agents. Like our analysis on $\epsilon$ for RL Max Loss, an optimal value of $c$ depends on the adversary’s goals and capabilities.
4.4 Utility Analysis
Each adversarial agent trains on images from the training dataset $\mathcal{D}$. To investigate a broader landscape of attack capabilities in black-box settings, we ask: how well do trained adversarial agents generalize to unseen data, and how does it compare to traditional black-box methods? We bifurcate the analysis of this question into two parts: (1) analyzing the adversarial samples crafted on an unseen dataset and (2) comparing the performance on unseen data of the RL attacks against known highly-performant black-box attacks on image classification. The first part will examine adversarial samples crafted after training on the training dataset $\mathcal{D}$ and testing dataset $\mathcal{D}^{\prime}$. The second part will compare the RL attacks against the black-box attack baselines over ASR, AQ, and average $\ell_{2}$-norm distortion. In Table 2, we take RL Max Loss and RL Min Norm agents after 1 200 PPO policy updates training on dataset $\mathcal{D}$ in CIFAR-10 and SVHN tasks and evaluate the trained agents on the respective training and testing datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$. Additionally, we evaluate the baselines on the testing datasets $\mathcal{D}^{\prime}$ in both tasks. This is to represent a comparison across attack methods where all have no experience attacking the same data. We report the ASR, AQ, and average $\ell_{2}$-norm distortion ($\ell_{2}$) across all methods and datasets.